-8(4+9x)=7(-2-11x)
Answer by ronalee(3)

The first thing you want to do is on the left side, distribute the -8 to the four and the 9x, and the problem becomes this:
-8(4+9x) =7(-2-11x)---> -32-72x= 7(-2-11x).
Next, you'll want to distribute the 7 on the right side to the -2 and the -11x, and the problem becomes this:
-32-72x=7(-2-11x)---> -32-72x= -14-77x
Now you can combine like terms; add the 32 to both sides, and add 77x to both sides:
-32(+32) - 72x(+77x) = -14(+32) - 77x(+77x).
And you will get:
5x= 18. Then divide both sides by 5 and x= 18/5